OpenStudy (anonymous):

What is the probability of getting all diamonds if 10 cards are selected without replacement

jimthompson5910 (jim_thompson5910):

There are 52 cards total There are 13 diamond cards. So on the first pick, the probability is 13/52 There are now 13 - 1 = 12 diamond cards out of a total of 52 - 1 = 51 cards So on the second pick, the probability is 12/51 There are now 12 - 1 = 11 diamond cards out of a total of 51 - 1 = 50 cards So on the third pick, the probability is 11/50 ....etc... this pattern keeps going... on the last pick (10th pick), there are only 4 diamonds left out of 43 cards total left over So the last probability is 4/43 Now multiply all the probabilities together to get (13/52)(12/51)(11/50)(10/49)(9/48)(8/47)(7/46)(6/45)(5/44)(4/43) 1037836800/57407703889536000 1/55314770 So the probability is 1/55314770